OpenStudy (anonymous):

Find the derivative of y = 14^sin(3x)

OpenStudy (cwrw238):

take logs ln y = sin(3x)* ln 14 1/y * dy/dx = ln14* 3 sin3x dy/dx = y * ln14 3sin3x = 14^sin(3x)* ln 14 * 3 sin (3x)

OpenStudy (anonymous):

:/ my quiz hw, says its wrong, idk what to do

OpenStudy (cwrw238):

what does it say the answer is?

OpenStudy (cwrw238):

oh right - sorry - its 3 cos 3x at the end

OpenStudy (anonymous):

Idk, I have one try left, you get three trys until its wrong

OpenStudy (cwrw238):

the derivative of sin 3x by chain rule is 3 cos3x not 3 sin3x

OpenStudy (anonymous):

so what would that make it 14^sin(3x)*ln14*3cos(3x)

OpenStudy (cwrw238):

yea - but ill check it out with wolfram first wait a sec

OpenStudy (anonymous):

okay (:

OpenStudy (anonymous):

it gave me a weird answer when i did it and i tried it and it said it was wrong

OpenStudy (cwrw238):

wolfram gets same as me but writes it in different order: 3 ln 14 14^sin(3x) cos (3x)

OpenStudy (cwrw238):

you can be sure thats correct

OpenStudy (cwrw238):

http://www.wolframalpha.com/input/?i=differentiate+14%5E%28sin+3x%29 - it uses log instead of ln but its same thing

OpenStudy (anonymous):

Yay, It's right, thank you so much!

OpenStudy (anonymous):

okay, I got another, What is the answer to this: calculate g'(−2), where g is the inverse of f(x) = 4x^3 - 2x

OpenStudy (anonymous):

or what do you get?

OpenStudy (cwrw238):

finding the inverse of that is beyond me - it would be a very complicated expression - unless there is a shortcut - thats beyond me too im afraid

OpenStudy (anonymous):

lol alright, thanks for the help though (: